OpenStudy (anonymous):

f(x)= 3/(x+6) g(x)=x^2-3 find(fg)(x)

OpenStudy (anonymous):

\[fg(x)=f(x)\times g(x)\] your job is to multiply \[\frac{3}{x+6}\times (x^2-3)\]

OpenStudy (anonymous):

Yes i know that but I dont know how solve the problem

OpenStudy (anonymous):

there i nothing really to do

OpenStudy (anonymous):

you can write \[\frac{3(x^2-3)}{x+6}\] i suppose

OpenStudy (anonymous):

thank you!